I need help!
solve each system of equations by elimination
5x - 4y = 23
7x + 8y = 5
Thanks

Answer by benni1013(206)

5x - 4y = 23
7x + 8y = 5
First thing you need to look at which variable is going to be easier to eliminate in this case it is the y. Multiply the top by two to get 10x-8y=46.
Adding the equations together this will ELIMINATE the y variable yielding.
17x=51 or x=3. Now plug this answer into one of the ORIGINAL equation and solving for y.